# Distribution estimation via Flow Matching with Lipschitz guarantees

### v1

@@ 7.4 Additional proofs of Section˜7 and helper results
$\mathbb{E}\Big{[}\big{(}Y_{i}^{2}-\mathbb{E}[Y_{i}^{2}]\big{)}^{k}\Big{]}$ $\lesssim\big{(}\frac{\sigma_{t}^{2}}{t^{2}+\sigma_{t}^{2}}\big{)}^{\frac{k}{2}}\leq\min(\frac{\sigma_{t}^{k}}{t^{k}},1).$
Now we consider the case that $f$ is of the form (45) or (47). First, we note that as shown in the proof of Theorem˜4.6, the distribution with density $q_{x,t}$ satisfies the Poincaré inequality with a constant
$\rho\geq e^{-2L}\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}.$
An immediate consequence of the Poincaré inequality Definition˜7.3 is the following bound for all smooth functions $g,h$
@@
We are going to use this to bound
$\operatorname{Cov}(f(Y)(\frac{\partial}{\partial t}\log(w_{x,t}(Y)))^{k-1},\frac{\partial}{\partial t}\log(w_{x,t}(Y))).$
Define
$g(y)\coloneqq f(y)(\frac{\partial}{\partial t}\log(w_{x,t}(y)))^{k-1}.$
Then
$\nabla g(y)=(\nabla f(y))(\frac{\partial}{\partial t}\log(w_{x,t}(y)))^{k-1}+f(y)(k-1)\nabla(\frac{\partial}{\partial t}\log(w_{x,t}(y)))^{k-2}(\nabla\frac{\partial}{\partial t}\log(w_{x,t}(y)))$
@@
$|\nabla g(y)|^{2}\lesssim|\nabla f(y)|^{2}|\frac{\partial}{\partial t}\log(w_{x,t}(y)))|^{2(k-1)}+|f(y)|^{2}(k-1)^{2}|(\frac{\partial}{\partial t}\log(w_{x,t}(y)))|^{2(k-2)}|\nabla\frac{\partial}{\partial t}\log(w_{x,t}(y))|^{2}$
@@
$\leq\frac{e^{2L}\operatorname{polylog}(\sigma_{\min})}{\int\exp(-\frac{|z|^{2}}{2}-\frac{|\frac{x-\sigma_{t}z}{t}|^{2}}{2})\;\mathrm{d}y}$
$\quad\cdot\int\operatorname{poly}_{2}(x,\frac{x-\sigma z}{t},t)\sigma_{t}^{-4(k-1)}|\sigma_{t}z|^{2(k-1)}|\frac{x-\sigma_{t}z}{t}|^{2(k-1)}\exp(-\frac{|z|^{2}}{2}-\frac{|\frac{x-\sigma_{t}z}{t}|^{2}}{2})\;\mathrm{d}y$
@@
where the last inequality stems from similar arguments as in the proof of Theorem˜4.6 and Theorem˜5.2 for higher order moments. The exact calculations are omitted for the sake of brevity. The same calculation for the other terms in (49), (50), (51) yield
@@
where we used that $x\in[-\log(n),\log(n)]^{d}$ to combine the worst case dependencies. For $h$, we obtain in similar manner
$\mathbb{E}[|\nabla h(y)|^{2}]\lesssim\sigma_{t}^{-4}\log(n)^{4}.$
Inserting these bounds in (48), we obtain
@@
$\frac{\partial^{k}}{\partial t^{k}}v^{j}_{t}(x)\lesssim\operatorname{polylog}(\sigma_{\min})\operatorname{polylog}(n)\sigma_{\min}^{-k-2}.$
This concludes the proof. ∎
7.4.1 Helper results
Lemma 7.10.
Let $\sigma_{t}=\sigma_{\min}^{t}$ with $\sigma_{\min}\in(0,1)$. Then for $t\in(0,1)$
$\frac{t}{t^{2}+\sigma_{t}^{2}}\leq\max(\log(\sigma_{\min}^{-1}),e^{2}).$
@@ 7.4.1 Helper results
First, let $t\geq\frac{1}{\log(\sigma_{\min}^{-1})}$. Then
$\frac{t}{t^{2}+\sigma_{t}^{2}}\leq\frac{1}{t}\leq\log(\sigma_{\min}^{-1}).$
If $t<\frac{1}{\log(\sigma_{\min}^{-1})}$, then
$\sigma_{\min}^{2t}=\exp(2\log(\sigma_{\min})t)\geq\exp(-2)$
and thus
$\frac{t}{t^{2}+\sigma_{t}^{2}}\leq\frac{t}{t^{2}+\exp(-2)}\leq\exp(2).$
Lemma 7.11.
@@
$\int\Big{|}\int_{0}^{1}\langle x-y,\nabla p^{*}(x+t(y-x)-z)\rangle\Big{|}\;\mathrm{d}z$ $\leq|x-y|\int\int_{0}^{1}|\nabla p^{*}(x+t(y-x)-z)|\;\mathrm{d}t\;\mathrm{d}z$
$=|x-y|\int_{0}^{1}\int|\nabla p^{*}(w)|\;\mathrm{d}w\;\mathrm{d}t.$
@@
$\int|\nabla p^{*}(w)|\;\mathrm{d}w$ $=\int|-w-\nabla a(w)|\frac{\exp(-\frac{-|w|^{2}}{2}-a(w))}{\int\exp(-\frac{-|w|^{2}}{2}-a(w))\;\mathrm{d}w}\;\mathrm{d}w$
$\leq\mathbb{E}_{W\sim p^{*}}[|W|]+\sqrt{d}L.$
@@
$\Big{|}\int\nabla f(z)p^{*}(x-z)\mathrm{d}z-\int\nabla f(z)p^{*}(y-z)\;\mathrm{d}z\Big{|}\leq d(e^{2L}+L)|x-y|.$

### v2

@@ Appendix A Additional proofs of Section˜6 and helper results
$\mathbb{E}\Big[\big(Y_{i}^{2}-\mathbb{E}[Y_{i}^{2}]\big)^{k}\Big]\lesssim\Big(\frac{\sigma_{t}^{2}}{t^{2}+\sigma_{t}^{2}}\Big)^{\frac{k}{2}}\leq\min\Big(\frac{\sigma_{t}^{k}}{t^{k}},1\Big).$
Now we consider the case that $f$ is of the form (43) or (45). First, we note that as shown in the proof of Proposition˜3.6, the distribution with density $q_{x,t}$ satisfies the Poincaré inequality with a constant $\rho\geq e^{-2L}\big(1+\frac{t^{2}}{\sigma_{t}^{2}}\big).$ An immediate consequence of the Poincaré inequality Definition˜6.3 is the following bound for all smooth functions $g,h$
@@
We are going to use this to bound $\operatorname{Cov}(f(Y)(\frac{\partial}{\partial t}\log(w_{x,t}(Y)))^{k-1},\frac{\partial}{\partial t}\log(w_{x,t}(Y))).$ Define $g(y)\coloneqq f(y)(\frac{\partial}{\partial t}\log(w_{x,t}(y)))^{k-1}.$ Then
$\nabla g(y)=(\nabla f(y))\Big(\frac{\partial}{\partial t}\log(w_{x,t}(y))\Big)^{k-1}+f(y)(k-1)\nabla\Big(\frac{\partial}{\partial t}\log(w_{x,t}(y))\Big)^{k-2}\Big(\nabla\frac{\partial}{\partial t}\log(w_{x,t}(y))\Big)$
@@
$|\nabla g(y)|^{2}\lesssim|\nabla f(y)|^{2}\Big|\frac{\partial}{\partial t}\log(w_{x,t}(y)))\Big|^{2(k-1)}+|f(y)|^{2}(k-1)^{2}\Big|\Big(\frac{\partial}{\partial t}\log(w_{x,t}(y))\Big)\Big|^{2(k-2)}\Big|\nabla\frac{\partial}{\partial t}\log(w_{x,t}(y))\Big|^{2}$
@@
$\leq\frac{e^{2L}\operatorname{polylog}(\sigma_{\min})}{\int\exp(-\frac{|z|^{2}}{2}-\frac{|\frac{x-\sigma_{t}z}{t}|^{2}}{2})\;\mathrm{d}y}\int\operatorname{poly}_{2}\Big(x,\frac{x-\sigma z}{t},t\Big)\sigma_{t}^{-4(k-1)}|\sigma_{t}z|^{2(k-1)}\Big|\frac{x-\sigma_{t}z}{t}\Big|^{2(k-1)}\exp\Big(-\frac{|z|^{2}}{2}-\frac{|\frac{x-\sigma_{t}z}{t}|^{2}}{2}\Big)\;\mathrm{d}y$
@@
where the last inequality stems from similar arguments as in the proof of Proposition˜3.6 and Theorem˜4.2 for higher order moments. The exact calculations are omitted for the sake of brevity. The same calculation for the other terms in (47), (48), (49) yield
@@
where we used that $x\in[-\log(n),\log(n)]^{d}$ to combine the worst case dependencies. For $h$, we obtain in similar manner $\mathbb{E}[|\nabla h(y)|^{2}]\lesssim\sigma_{t}^{-4}\log(n)^{4}.$ Inserting these bounds in (46), we obtain
@@
$\frac{\partial^{k}}{\partial t^{k}}v^{j}_{t}(x)\lesssim\operatorname{polylog}(\sigma_{\min})\operatorname{polylog}(n)\sigma_{\min}^{-k-2}.\qed$
A.0.1 Helper results
Lemma A.2.
[…]
@@ A.0.1 Helper results
First, let $t\geq\frac{1}{\log(\sigma_{\min}^{-1})}$. Then $\frac{t}{t^{2}+\sigma_{t}^{2}}\leq\frac{1}{t}\leq\log(\sigma_{\min}^{-1}).$ If $t<\frac{1}{\log(\sigma_{\min}^{-1})}$, then $\sigma_{\min}^{2t}=\exp(2\log(\sigma_{\min})t)\geq\exp(-2)$ and thus $\frac{t}{t^{2}+\sigma_{t}^{2}}\leq\frac{t}{t^{2}+\exp(-2)}\leq\exp(2).$
Lemma A.3.
@@
$\int\Big|\int_{0}^{1}\langle x-y,\nabla p^{*}(x+t(y-x)-z)\rangle\Big|\;\mathrm{d}z$ $\leq|x-y|\int_{0}^{1}\int|\nabla p^{*}(w)|\;\mathrm{d}w\;\mathrm{d}t.$
@@
$\int|\nabla p^{*}(w)|\;\mathrm{d}w=\int|-w-\nabla a(w)|\frac{\exp(-\frac{-|w|^{2}}{2}-a(w))}{\int\exp(-\frac{-|w|^{2}}{2}-a(w))\;\mathrm{d}w}\;\mathrm{d}w\leq\mathbb{E}_{W\sim p^{*}}[|W|]+\sqrt{d}L.$
@@
$\Big|\int\nabla f(z)p^{*}(x-z)\mathrm{d}z-\int\nabla f(z)p^{*}(y-z)\;\mathrm{d}z\Big|\leq d(e^{2L}+L)|x-y|.\qed$